OpenStudy (anonymous):

How do you find the asymptotes of a trigonometric function? Say the function is f(x)=tan2x and the asymptotes are from x=0 to x=pi. I know how to solve for asymptotes, I just don't know what to do when they give me a "range"

OpenStudy (jdoe0001):

well... a range just means " what vertical asymptotes occur between here and here"

OpenStudy (jdoe0001):

and tangent function has infinite vertical asymptotes, since its period is based on it

OpenStudy (anonymous):

infinite? i don't see that as a solution

OpenStudy (anonymous):

the question asks *where the asymptotes are

OpenStudy (jdoe0001):

heheh... I said, that tangent itself, has infinite since it runs infintely

OpenStudy (jdoe0001):

well.... the asymptotes will lie where the period for tangent starts and end so for tan(2x) for example the period for that function will be of \(\bf tan({\color{brown}{ 2}}x)\qquad period\implies \cfrac{\pi}{{\color{brown}{ 2}}}\)

OpenStudy (jdoe0001):

period of a trig function, is the "multiplier" of the angle, dividing by the "regular period" of it tangent has a regular period of \(\pi\)

OpenStudy (anonymous):

mmm.. I know what the period is. but how would i use it to find where the asymptotes are?? do i add pi/2b to it for one of them?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

so i get 5pi/4 for the first asymptote... but is it at x=0??

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

the period for this one is \(\bf tan({\color{brown}{ 2}}x)\qquad period\implies \cfrac{\pi}{{\color{brown}{ 2}}}\) not \(\bf \cfrac{\pi}{4}\)

OpenStudy (jdoe0001):

so, there's one at 0 then \(\bf\cfrac{\pi}{2}, \cfrac{\pi}{2}+\cfrac{\pi}{2}\)

OpenStudy (anonymous):

so i add pi/2 with pi/4?

OpenStudy (anonymous):

because pi/2b is pi/4

OpenStudy (jdoe0001):

anyhow you'd go from one period to the next to the next at every period stop, the x-coordinate is the vertical asymptote

OpenStudy (anonymous):

i don't exactly follow. can i just show you the possible solutions to them? and you can explain to me which one is correct? working backwards?

OpenStudy (anonymous):

1: x=0, x=pi/4 2: x=pi/4, x=3pi/4 3. x=pi/2, x=3pi/2

OpenStudy (anonymous):

... better to graph the function to see where the asymptotes are...